Grade 11Algebra

p,q,r are distinct numbers in geometric progression and P + x, q + x, r+x , are in harmonic progression then x is equal to

Jwalin Raval
9 Years agoGrade 11

1 Answer

Jitender K Yadav
9 Years ago
x=q
This can be solved as under:-
p/q=q/r, which implies that q2=pr.        -(1)
(p+x), (q+x), (r+x) are in HP,
Thus, \tfrac{2}{q+x}=\tfrac{1}{p+x}+\tfrac{1}{r+x}
2(p+x)(r+x)=(q+x)(p+r+2x)
or (2pr+2px+2rx+2x2)=(pq+rq+2qx+px+rx+2x2)
which can be written as:-
x(2p+2r-p-r-2q)=(pq+rq-2pr)           -(2)
LHS1=x(p+r-2q)                             -(2L)
RHS1=[(p+r)q-2pr]=[(p+r)q-2q2]                         [From eq. 1]
RHS1=(p+r-2q)q=q(LHS1)/x            -(2R)            [From eq. 2L]
This implies that:-
LHS1=q(LHS1)/x                                                [From eqs. 2, 2L, 2R]
x=q